OpenStudy (anonymous):

The ages of Max and his father add up to 35 years. Max’s father’s age is the same as five more than the square of Max’s age. a) Write a system of equations to represent this situation. Define your variables. b) Solve the system graphically. Are all possible solutions meaningful? Explain. c) How old are Max and his father?

OpenStudy (anonymous):

How do i make the system if equation?

OpenStudy (crashonce):

ok so let max's age be A and his father's age be B

OpenStudy (anonymous):

Okay

OpenStudy (crashonce):

therefore we get two equations: A+B=35 and B=A^2 =5 understand?

OpenStudy (anonymous):

can we use x and y please it gets confusing if i add something else

OpenStudy (crashonce):

sure: x+y=35 and x=y^2 =5 where x is the age of Max and y is the age of his father now thats part a done

OpenStudy (crashonce):

sorry x=y^2+5

OpenStudy (anonymous):

x=35-y then i add that into the other equation? so: 35-y=y^2+5

OpenStudy (crashonce):

yup

OpenStudy (anonymous):

okay now theres a y and a y^2 .... i should transfer the -y to the other side

OpenStudy (crashonce):

sorry x is the age of the father

OpenStudy (crashonce):

but yes you do

OpenStudy (crashonce):

now solve the equation

OpenStudy (anonymous):

ugh confused

OpenStudy (anonymous):

theres no x left though

OpenStudy (anonymous):

so are we solving for max now?

OpenStudy (crashonce):

so after moving everything to one side you should get: \[y^2 + y - 30 = 0\]

OpenStudy (crashonce):

yes solving for Max

OpenStudy (anonymous):

I agree

OpenStudy (crashonce):

now by solving the equation, you will find that max has two ages: 5 and -6. Because you cannot be -6 years old, Max is 5 Therefore his father is 35-5 = 30yrs old

OpenStudy (anonymous):

okay now we factor _+_=1 _x_=-30 right?

OpenStudy (anonymous):

okay can you explain how you did that?

OpenStudy (crashonce):

yes so it is =6 and -5 because +6-5=1 and 6 x -5=-30

OpenStudy (anonymous):

i should've spent more days factoring :(

OpenStudy (crashonce):

it gets very easy when you know how to do it

OpenStudy (anonymous):

I know how but the problem is i don't see it right away :( it takes me too long. thank you, now im going to revise this and try to solve it using the graph cause it tells me to do so

OpenStudy (crashonce):

no worries

OpenStudy (anonymous):

How do I sketch this?